Circular No. 5805

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET NEUJMIN 3 (1993j) J. V. Scotti, Lunar and Planetary Laboratory, reports his recovery of

this comet with the 0.9-m Spacewatch telescope at Kitt Peak. The comet is essentially of stellar appearance. The indicated correction to the prediction on MPC 16381 (ephemeris on MPC 21285) is ΔT = +0.03 day.

1993 UT  α_{2000}  δ_{2000}  m_{1}
May 25.34975  15^{h}03^{m}29.^{s}83  −10^{o}32′37′′.9  21.0
25.38166  15 03 28.38  −10 32 33.5  21.2
26.24980  15 02 52.45  −10 30 10.3  22.1
26.29348  15 02 50.59  −10 30 01.9  21.7
26.31215  15 02 49.77  −10 29 59.2  21.9

YY HERCULIS This Z And-type variable has brightened for the first time since mid-

1981, as indicated by the following visual magnitude estimates: Apr. 24.2 UT, 13.1 (J. E. Bortle, Stormville, NY); 26.02, 13.2 (J. Speil, Walbrzych, Poland); 29.3, 13.2 (Bortle); May 9.1, 12.9 (Bortle); 12.95, 12.9 (Speil); 15.1, 12.7 (Bortle); 16.98, 12.6 (Speil); 18.96, 12.6 (Speil); 22.1, 12.5 (Bortle); 23.1, 12.4 (Bortle); 26.1, 12.1 (Bortle). J. A. Mattei, AAVSO, notes that YY Her had faded to m_{v} = 14.3 in 1992 Oct.-Nov.

PSR 2334+61 W. Becker and J. Trümper, Max-Planck-Institut für Extraterrestrische

Physik; and H. Ogelman, University of Wisconsin, report: “We have detected the pulsar PSR 2334+61 (dynamic age ∼ 41 000 yr) in the energy range 0.1-2.4 keV with ROSAT. The source was observed with the PSPC from 1992 July 27 to Aug. 4, for a total exposure time of 8350 s. The net count rate (including background, vignetting, and deadtime corrections) was 0.0018 ± 0.0005 count/s. The low x-ray flux precluded any detailed spectral analysis and modulation testing at the 0.495-s radio period. Assuming that the detected x-ray emission is of thermal origin, the canonical neutron-star parameters (R = 10 km, 1.4 M_{⊙}) and the parameters given in the Princeton Pulsar Catalog (d = 2.46 kpc, DM = 57.6 pc cm^{−3}), we find a surface temperature of log(T_{s},K) = 6.0327. The result is in good agreement with the predictions of standard neutron-star cooling theories.”

1993 May 27 Daniel W. E. Green